You found out that now you are going to receive payments of $8,500 for the next 13 years. You will receive these payments at the beginning of each year. The annual interest rate will remain constant at 11%. What is the present value of these payments? (Note: Round your answer to the nearest whole dollar.)
Answer #1

Payments received at the beginning is called Annuity due.

When the payments are received at the beginning of a year, it is called as annuity due
Present Value of annuity due is calculated as:
P*[(1-(1+r)^(-n)) / r] * (1+r)

P refers to the periodic payments = 8500
r refers to rate per period = 11%
n refers to the number of periods = 13
Present Value of annuity =8500*((1-(1+0.11)^(-13))/0.11*(1+0.11))
=8500*((1-0.257514256)/0.11*(1.11))
=8500*(0.742485744/0.11*(1.11))
=63685.027 or $63685 (Rounded to nearest whole dollar)